OpenStudy (anonymous):

The speed of a stream is 4 mph. If a boat travels 44 miles downstream in the same time that it takes to travel 22 mile upstream, what is speed of the boat in still water?

OpenStudy (lgbasallote):

let x = speed of boat y = current downstream = x + y upstream = x -y t = distance/speed 44/x + 4 = 22/x - 4 44x - 176 = 22x + 88 22x = 264 x = 12 mph